An automobile cooling system contains 36 liters
of coolant that is 14% antifreeze. How many liters of the mixture must be
removed so that, when it is replaced with pure antifreeze, a mixture of
37% antifreeze will result?

Answer by ankor@dixie-net.com(22740)

An automobile cooling system contains 36 liters of coolant that is 14% antifreeze.
How many liters of the mixture must be removed so that, when it is replaced with pure antifreeze, a mixture of 37% antifreeze will result?
:
let a = amt of pure antifreeze to be added, also amt to be removed
:
.14(36-a) + a = .37(36)
5.04 - .14a + a = 13.32
-.14a + a = 13.32 - 5.04
.86a = 8.28
a ~ 9.63 liters removed then that same amt of pure antifreeze is added